# PEAR: Equal Area Weather Forecasting on the Sphere

### v1

@@ Abstract
Machine learning methods for global medium-range weather forecasting have recently received immense attention. Following the publication of the Pangu Weather model, the first deep learning model to outperform traditional numerical simulations of the atmosphere, numerous models have been published in this domain, building on Pangu’s success. However, all of these models operate on input data and produce predictions on the Driscoll–Healy discretization of the sphere which suffers from a much finer grid at the poles than around the equator. In contrast, in the Hierarchical Equal Area iso-Latitude Pixelization (HEALPix) of the sphere, each pixel covers the same surface area, removing unphysical biases. Motivated by a growing support for this grid in meteorology and climate sciences, we propose to perform weather forecasting with deep learning models which natively operate on the HEALPix grid. To this end, we introduce Pangu Equal ARea (PEAR), a transformer-based weather forecasting model which operates directly on HEALPix-features and outperforms the corresponding model on Driscoll–Healy without any computational overhead. Code will be made available.
@@ 1 Introduction
Following the publication of the landmark model Pangu-Weather [bi2023], machine learning has been rapidly transforming the field of weather forecasting, demonstrating remarkable potential for predicting key meteorological variables such as temperature, pressure and winds [lam2023, chen2023, chen2023a, andrychowicz2023a, NEURIPS2024_7f19b99e]. This progress has been driven by the widespread use of the ERA5 reanalysis dataset [hersbach2020era5], a comprehensive open access global climate data record produced by the European Centre for Medium-Range Weather Forecasts (ECMWF).
@@
ERA5 provides hourly estimates of atmospheric, land, and oceanic variables on a global scale, discretized using the regular latitude-longitude grid or its spherical harmonic counterpart, the Driscoll–Healy grid [driscoll1994]. While this grid enables spectral representation and fast spherical transforms, it introduces challenges when used in physics-based ML models. Specifically, the Driscoll–Healy grid exhibits non-uniform resolution across latitudes and does not preserve the symmetries of the sphere, which can lead to artefacts and unphysical inductive biases in learning-based models.
@@
To address these limitations, alternative spherical representations have been explored. One notable example is HEALPix (Hierarchical Equal Area isoLatitude Pixelation), originally developed in astrophysics for uniform pixelization of the celestial sphere  [gorski1998analysis]. HEALPix offers equal-area pixel distributions and hierarchical resolution capabilities, making it well-suited for learning tasks on spherical domains. Recent efforts have begun to investigate the use of HEALPix for resampling next-generation weather and climate data [karlbauer2024, ramavajjala2024]. HEALPix has also been used in other machine learning contexts where spherical data is natural [Perraudin_2019, carlsson2023].
@@
The ECMWF has also acknowledged the potential of HEALPix in its Destination Earth (DestinE) initiative [HOFFMANN2023100394], which aims to develop a digital twin of the Earth for improved climate and weather predictions. By employing HEALPix, DestinE seeks to achieve high-resolution, globally consistent datasets that can better inform decision-making processes related to climate change adaptation and disaster risk management. Furthermore, ECMWF’s transition from GRIB1 to GRIB2 data formats supports the integration of advanced grid systems like HEALPix [grib1togrib2], enabling more detailed and efficient data representation essential for next-generation forecasting models.
@@
Motivated by these developments, we propose PEAR (Pangu Equal Area), the first machine learning weather forecasting model that operates entirely on the HEALPix grid. Unlike previous approaches that project onto planar grids or only use HEALPix for preprocessing, PEAR natively represents inputs, internal features, and outputs on the spherical HEALPix domain, enabling consistent resolution across the globe and better alignment with the underlying physical symmetries. We use ERA5 data as the forecasting target, resampled to HEALPix, and evaluate our model’s performance on key atmospheric variables across various resolutions.
@@
  • Motivated by an increased adoption of the HEALPix grid in next generation weather and climate data, we propose to use the spherical HEALPix grid as the native grid for machine learning weather predictions. This approach eliminates unphysical biases in the sampling of the sphere affecting all standard Driscoll–Healy based weather prediction models and removes the need for spatial weights in the loss and evaluation metrics.
  • We introduce PEAR: a baseline model for weather prediction on HEALPix using a volumetric transformer architecture which operates at no computational overhead compared to an equivalent model on the traditional Driscoll–Healy grid.
  • To demonstrate the advantage of using a native HEALPix model, we show that PEAR’s HEALPix-predictions outperform those produced by the same architecture operating on the Driscoll–Healy grid. We evaluate all models for lead times of up to ten days and show that this advantage persists for longer forecasting horizons, even though the Driscoll–Healy-based model has more than twice as many parameters. PEAR also outperforms Pangu-Large with almost eight times as many parameters at a forecast horizon of 5 days and beyond.
@@ 2 Related work
The field of machine-learning weather forecasting has received tremendous attention over the last years both for medium-range weather forecasting [shu2024] and extreme weather prediction [olivetti2024]. The first model whose performance on global medium-range forecasts surpassed that of numerical models was Pangu-Weather [bi2023] which is based on a volumetric version of the SWIN transformer [liu2021a]. Since then, a number of models have been published which improved upon this baseline, such as GraphCast [lam2023], FuXi [chen2023], FengWu [chen2023a], NetMet-3 [andrychowicz2023a], Stormer [NEURIPS2024_7f19b99e] or the ECMWF’s data-driven forecasting system [lang2024]. Fourier neural operators have been used in a number of models in this domain [kurth2023, bonev2023] as well. In order to take the curvature of the sphere into account, one model [PhysicsInspiredAdaptionstoLowParameterNeuralNetworkWeatherForecastSystems] used Spherenet. Similarly, the forecasting model CirT [liu2025cirt] is based on a circular transformer which takes the azimuthal circularity of the sphere into account. Probabilistic weather forecasting models allow for uncertainty estimation of the generated predictions [oskarsson2024]. The model GenCast creates an ensemble of stochastic predictions which outperforms the top operational medium-range weather forecast in the world, ENS, the ensemble forecast of the European Centre for Medium-Range Weather Forecasts [price2024]. The consideration of physical conservation laws in the training process can improve these data-driven weather prediction models [sha2025]. Closely related to machine-learning weather forecasting systems are neural network models which predict the climate [nguyen2023a, watt-meyer2023] or general-purpose models for the earth system [bodnar2024]. Contributions to the training process include careful ablations of various aspects of the architecture [AnalyzingandExploringTrainingRecipesforLargeScaleTransformerBasedWeatherPrediction] and a training platform for deep-learning based weather prediction models [schreck2024]. WeatherBench 2 [rasp2024] provides a well-established benchmark for machine-learning weather prediction models. Deep neural networks have also been applied in the post-processing of global weather forecasts [bouallegue2024]. A hybrid model of machine-learning components and a differentiable solver for circulation models was proposed in [kochkov2024].
@@
The aforementioned models have in common that they use the naive Driscoll-Healy discretization [driscoll1994] of the sphere. The resulting grid is considerably denser towards the poles than around the equator. In contrast, the HEALPix grid is uniform over the entire sphere. The convolutional DLWP-HPX model [karlbauer2024] uses representations on the HEALPix grid. HEAL-ViT [ramavajjala2024] is a vision-transformer based model which operates on the HEALPix grid and uses a learnable encoder and decoder to map the features from the Driscoll-Healy grid to HEALPix and back. In contrast, our model operates entirely on HEALPix, in line with our proposal to use the HEALPix grid as the physically appropriate grid for weather forecasting.
@@ 3.2 ERA5 on HEALPix
The ERA5 dataset [hersbach2020era5] contains the global atmospheric state in terms of a number of hydrodynamical quantities, discretized on a Driscol-Healy gridding on the sphere, over multiple vertical slices and hourly in time. Following Pangu-Weather [bi2023], we distinguish between surface variables (temperature, wind velocity, pressure, humidity) and upper variables (temperature, wind velocity, humidity, pressure).
The raw data provided by ECWMF has an angular gridding of $0.25^{\circ}$ resulting in a spatial $(\mathrm{lon},\mathrm{lat})$ resolution of $(1440,721)$. Since the Driscoll-Healy grid cells are not of equal area, the effective spatial resolution on the sphere is not constant. The lowest spatial resolution can be found on the equator, where each cell has an angular area of $1.9\cdot 10^{-5}\mathrm{rad}^{2}$. We target a HEALPix grid with $n_{\mathrm{side}}=64$, where all of the $12\cdot 64^{2}$ pixels cover an angular area of $2.6\cdot 10^{-4}$, so that the ERA5 data has a higher spatial resolution everywhere on the globe and saturates the HEALPix pixel density.
@@ 3.3 ERA5-lite
Table 1: Tensor shapes of a single sample regridded to the HEALPix grid with $n_{\mathrm{side}}=64$. The tensor shape for the upper variables includes the subset of 13 pressure levels from the 37 available in the ERA5 dataset.
Tensor Shape
surface $(49152,4)$
upper $(49152,13,5)$
@@ 4 PEAR: Pangu Equal Area
The architecture is constructed using a combination of 5 main layer types: patch embedding, windowed attention with alternating shifting, downsampling, upsampling and patch recovery. See Figure 4 for a schematic overview of the architecture, and Table 2 for details on the layer parameters.
@@ 4.1 Patch embedding
The initial patch embedding uses a 1d convolution with kernel size 16 and stride 16 for the surface variables, and a 2d convolution with kernel size and stride $(16,2)$ for the upper variables. This corresponds to a patch size of $4\times 4$ in the Driscoll-Healy grid. Both convolutions output 48 channels. The patch embedded tensors have shape $(\nicefrac{{3}}{{4}}n_{\mathrm{nside}}^{2},48)$ and $(\nicefrac{{3}}{{4}}n_{\mathrm{nside}}^{2},7,48)$. At this point the patch embedded surface variables are concatenated to the patch embedded upper variables, resulting in a single tensor of shape $(\nicefrac{{3}}{{4}}n_{\mathrm{nside}}^{2},8,48)$.
@@ 4.2 Windowed attention
Figure 2: Patches (green) and windows (red) at the first attention layer. Each green patch contains four pixels of the input resolution $n_{\mathrm{side}}=64$ on the HEALPix grid, and each window contains 64 patches at this level. The window also include one vertical level above, not shown in the figure.
We partition the tensors into windows using the nested structure of the HEALPix grid. The vertical direction is also partitioned into neighbouring levels. In terms of the latent tensor shape $(\nicefrac{{3}}{{4}}n_{\mathrm{nside}}^{2},8,48)$, partitioning using a window size of $(W_{\mathrm{hp}},W_{d})$ gives a windowed tensor of shape $(\frac{\nicefrac{{3}}{{4}}n_{\mathrm{nside}}^{2}}{W_{\mathrm{hp}}}\frac{8}{W_{\mathrm{d}}},W_{\mathrm{hp}}W_{d},48)$. Attention is now performed over the embeddings of the $W_{\mathrm{hp}}W_{d}$ voxels in each window. Note that the nested index structure of the HEALPix grid naturally supports the window partitioning in terms of contiguous memory. See Figure 2 for an illustration of the patches and first level windows corresponding to the tensor structure at the first attention layer.
In contrast to Pangu [bi2023], we use a simplified learned relative positional embedding. Because of the equal area grid cells in the HEALPix grid we can share relative positional embeddings between windows, and thus simply learn a relative positional embedding tensor $B$ of shape $(1,N_{\mathrm{heads}},\left(W_{d}W_{\mathrm{hp}}\right)^{2})$. The final attention computation is thus
$\mathrm{Att}(Q,K,V)=\mathrm{SoftMax}\left(\frac{QK^{\top}}{\sqrt{d}}+B\right)V,$ (1)
where $d$ is the embedding dimension. The simplified positional embedding accounts for most of the parameter savings compared to Pangu in Table 3.
@@ 4.4 Patch recovery
Table 2: PEAR architecture overview. The windowed attention blocks contain multiple copies, indicated by the depth column, of the multi-head attention layer and layer normalization. The tensor shape column shows the shape of the output tensor from the corresponding layer. Both the input and final patch recovery layer use two separate tensors for the surface and upper variables.
Nr Layer/Block Depth Tensor shape Attention heads
1 Patch embed 1 $(\nicefrac{{3}}{{4}}n_{\mathrm{side}}^{2},8,48)$
2 Windowed attention 2 $(\nicefrac{{3}}{{4}}n_{\mathrm{side}}^{2},8,48)$ 6
3 Downsample 1 $(\nicefrac{{3}}{{16}}n_{\mathrm{side}}^{2},8,96)$
4 Windowed attention 12 $(\nicefrac{{3}}{{16}}n_{\mathrm{side}}^{2},8,96)$ 12
5 Upsample 1 $(\nicefrac{{3}}{{4}}n_{\mathrm{side}}^{2},8,48)$
6 Windowed attention 2 $(\nicefrac{{3}}{{4}}n_{\mathrm{side}}^{2},8,48)$ 6
Concatenate 6 & 2 1 $(\nicefrac{{3}}{{4}}n_{\mathrm{side}}^{2},8,96)$
@@ 5 Experiments
We evaluate PEAR against Pangu [zhuoqun2025] using the ERA5-lite subset described in Section 3.3. All experiments are carried out using $n_{\mathrm{side}}=64$ to saturate the angular resolution of the HEALPix grid using the available ERA5 data on the Driscoll-Healy grid. We train all models using L1 loss, with the loss contribution from the surface variables weighted by $\frac{1}{4}$. We use the AdamW optimizer with weight decay $3\times 10^{-6}$ and learning rate $5\times 10^{-4}$. On a single A100, PEAR converges in 20 hours and Pangu in 40 hours.
To evaluate the medium-term forecasting ability of PEAR, we perform iterated model inference up to 10 times, resulting in forward time predictions of up to 10 days. At each lead time we calculate the average RMSE and anomaly correlation coefficient (ACC) [van2017probability] of all variables over the globe according to
where $\Delta y$ is the difference between the predictions and the climatology average. To evaluate baseline-predictions on Driscoll–Healy, we apply the latitude weighting used in prior work [bi2023]. The equal area grid cells of HEALPix make this reweighting redundant for PEAR. The ACC measures the correlation between deviations from the climatology mean of predicted and ground truth forecasts, with a value of 1 indicating perfect agreement [hersbach2020era5]. The climatology average is subtracted to factor out seasonal variations that would otherwise improve the raw correlation between predictions and ground truth data. To calculate the climatology mean, we average each day of the year over the 11 years in the ERA5-lite dataset for every variable.
Pangu and Pangu-Large provide baselines against the same type of architecture where the Driscoll-Healy grid is used for input, latent and output tensors. Since the overall architectural blocks are the same, we can match the depth, number of attention heads and embedding dimensions for each block. Pangu and PEAR share the same architecture hyperparameters, whereas Pangu-Large has 4 times larger embedding dimension. For Pangu and Pangu-Large we use a Driscoll-Healy grid with $n_{\mathrm{lon}}=314$ and $n_{\mathrm{lat}}=157$, resulting in a relative difference of the total number of pixels of $0.3\%$, or 146 pixels.
@@
Table 4 lists the average ACC and RMSE for three different prediction lead times. PEAR consistently outperforms Pangu at longer lead times, and is often better than PanguLarge with 4 times larger embedding dimension.
See Appendix A for ACC and RMSE over individual pressure levels, as well as spatially resolved RMSE.

### v2

@@ Abstract
Artificial intelligence is rapidly reshaping the natural sciences, with weather forecasting emerging as a flagship AI4Science application where machine learning models can now rival and even surpass traditional numerical simulations. Following the success of the landmark models Pangu Weather and Graphcast, outperforming traditional numerical methods for global medium-range forecasting, many novel data-driven methods have emerged. A common limitation shared by many of these models is their reliance on an equiangular discretization of the sphere which suffers from a much finer grid at the poles than around the equator. In contrast, in the Hierarchical Equal Area iso-Latitude Pixelization (HEALPix) of the sphere, each pixel covers the same surface area, removing unphysical biases. Motivated by a growing support for this grid in meteorology and climate sciences, we propose to perform weather forecasting with deep learning models which natively operate on the HEALPix grid. To this end, we introduce Pangu Equal ARea (PEAR), a transformer-based weather forecasting model which operates directly on HEALPix-features and outperforms the corresponding model on an equiangular grid without any computational overhead.
https://github.com/hlinander/PEAR-Weather
@@ 1 Introduction
The convergence of artificial intelligence and the natural sciences—often referred to as AI4Science—is transforming how complex physical systems are studied and predicted. In Earth system science, machine learning approaches have achieved performance on par with or exceeding that of state-of-the-art numerical models [yu2024deep, liu2024foundation, zhao2024comprehensive]. Weather forecasting stands at the forefront of this shift [lam2023, chen2023, chen2023a, andrychowicz2023a, NEURIPS2024_7f19b99e], driven by advances in architectures that can capture multiscale atmospheric dynamics directly from data, while offering orders of magnitude speedups compared to traditional numerical integration methods. Yet, most existing AI-based forecasting systems inherit discretizations from the numerical models they aim to complement or replace, such as the equiangular latitude-longitude and Gaussian grids, which imposes non-uniform resolution and unphysical biases. In this paper we explore the Hierarchical Equal Area iso-Latitude Pixelization (HEALPix) as a native grid for learning-based weather forecasting, equal area modeling across the globe.
@@
Following the publication of the landmark models Pangu-Weather [bi2023] and Graphcast [lam2023], the recent emergence of performant data-driven models for weather forecasting has been driven by the widespread use of the ERA5 reanalysis dataset [hersbach2020era5], a comprehensive open access global climate data record produced by the European Center for Medium-Range Weather Forecasts (ECMWF).
@@
ERA5 provides hourly estimates of atmospheric, land, and oceanic variables on a global scale, discretized using the regular equiangular latitude-longitude grid or a reduced Gaussian grid. Both of these grids introduce challenges when used in physics-based ML models. Specifically, both grids exhibit non-uniform resolution across latitudes, which can lead to artifacts and unphysical inductive biases in learning-based models.
@@
To address these limitations, alternative spherical representations have been explored. One notable example is HEALPix, originally developed in astrophysics for uniform pixelization of the celestial sphere  [gorski1998analysis]. HEALPix offers equal-area pixel distributions and hierarchical resolution capabilities, making it well-suited for learning tasks on spherical domains. Recent efforts have begun to investigate the use of HEALPix for resampling next-generation weather and climate data [karlbauer2024, ramavajjala2024]. HEALPix has also been used in other machine learning contexts where spherical data is natural [Perraudin_2019, carlsson2023].
@@
The ECMWF are also targetting the HEALPix grid in its Destination Earth (DestinE) initiative [egusphere-2025-2198, HOFFMANN2023100394], which aims to develop a digital twin of the Earth for improved climate and weather predictions. By employing HEALPix, DestinE seeks to achieve high-resolution, globally consistent datasets that can better inform decision-making processes related to climate change adaptation and disaster risk management. Furthermore, ECMWF’s transition from GRIB1 to GRIB2 data formats supports the integration of advanced grid systems like HEALPix [grib1togrib2], enabling more detailed and efficient data representation essential for next-generation forecasting models.
@@
Motivated by these developments, we propose PEAR (Pangu Equal Area), the first machine learning weather forecasting model that operates entirely on the HEALPix grid. Unlike previous approaches that project onto planar grids or only use HEALPix for preprocessing, PEAR natively represents inputs, internal features, and outputs on the spherical HEALPix domain, enabling consistent resolution across the globe and better alignment with the underlying physical symmetries. We use ERA5 data, resampled to HEALPix, as the forecasting target and evaluate our model’s performance on key atmospheric variables across various resolutions.
@@
  • Motivated by an increased adoption of the HEALPix grid in next generation weather and climate data, we propose to use the spherical HEALPix grid as the native grid for machine learning weather predictions. This approach eliminates unphysical biases in the sampling of the sphere affecting standard equiangular based weather prediction models and removes the need for spatial weights in the loss and evaluation metrics.
  • We introduce PEAR: a baseline model for neural weather simulation on HEALPix using a volumetric transformer architecture which operates at no computational overhead compared to an equivalent model on the traditional equiangular grid.
  • To demonstrate the advantage of using a native HEALPix model, we show that PEAR’s HEALPix-predictions outperform those produced by the same architecture operating on an equiangular grid. We evaluate all models for lead times of up to ten days and show that this advantage persists for longer forecasting horizons, even though the equiangular-based model has more than twice as many parameters. PEAR also outperforms Pangu-Large with almost eight times as many parameters at a forecast horizon of 5 days and beyond.
@@ 2 Related work
The field of machine-learning weather forecasting has received tremendous attention over the last years both for medium-range weather forecasting [shu2024] and extreme weather prediction [olivetti2024]. The first model whose performance on global medium-range forecasts surpassed that of numerical models was Pangu-Weather [bi2023] which is based on a volumetric version of the SWIN transformer [liu2021a]. Since then, a number of models have been published which improved upon this baseline, such as GraphCast [lam2023], FuXi [chen2023], FengWu [chen2023a], NetMet-3 [andrychowicz2023a], Stormer [NEURIPS2024_7f19b99e] or the ECMWF’s data-driven forecasting system [lang2024]. Fourier neural operators have been used in a number of models in this domain [kurth2023, bonev2023] as well. In order to take the curvature of the sphere into account, one model [PhysicsInspiredAdaptionstoLowParameterNeuralNetworkWeatherForecastSystems] used Spherenet. Similarly, the forecasting model CirT [liu2025cirt] is based on a circular transformer which takes the azimuthal circularity of the sphere into account. Probabilistic weather forecasting models allow for uncertainty estimation of the generated predictions [oskarsson2024]. The model GenCast creates an ensemble of stochastic predictions which outperforms the top operational medium-range weather forecast in the world, ENS, the ensemble forecast of the European Centre for Medium-Range Weather Forecasts [price2024]. The consideration of physical conservation laws in the training process can improve these data-driven weather prediction models [sha2025]. Closely related to machine-learning weather forecasting systems are neural network models which predict the climate [nguyen2023a, watt-meyer2023] or general-purpose models for the earth system [bodnar2024]. Contributions to the training setup include careful ablations of various aspects of the architecture [AnalyzingandExploringTrainingRecipesforLargeScaleTransformerBasedWeatherPrediction] and a training platform for deep-learning based weather prediction models [schreck2024]. WeatherBench 2 [rasp2024] provides a well-established benchmark for machine-learning weather prediction models. Deep neural networks have also been applied in the post-processing of global weather forecasts [bouallegue2024]. A hybrid model of machine-learning components and a differentiable solver for circulation models was proposed in [kochkov2024].
@@
The aforementioned models have in common that they use the naive equiangular discretization of the sphere as input. The resulting grid is considerably denser towards the poles than around the equator. In contrast, the HEALPix grid is uniform over the entire sphere. The convolutional DLWP-HPX model [karlbauer2024] uses representations on the HEALPix grid. HEAL-ViT [ramavajjala2024] is a vision-transformer based model which operates on the HEALPix grid and uses a learnable encoder and decoder to map the features from the equiangular grid to HEALPix and back. In contrast, our model operates entirely on HEALPix, in line with our proposal to use the HEALPix grid as the physically appropriate grid for weather forecasting.
@@ 3.2 ERA5 on HEALPix
The ERA5 dataset [hersbach2020era5] contains the global atmospheric state in terms of a number of hydrodynamical quantities, discretized on an equiangular gridding on the sphere, over multiple vertical slices and hourly in time. Following Pangu-Weather [bi2023], we distinguish between surface variables (temperature, wind velocity, pressure, humidity) and upper variables (temperature, wind velocity, humidity, pressure).
The raw data provided by ECWMF has an angular gridding of $0.25^{\circ}$ resulting in a spatial $(\mathrm{lon},\mathrm{lat})$ resolution of $(1440,721)$. Since the equiangular grid cells are not of equal area, the effective spatial resolution on the sphere is not constant. The lowest spatial resolution can be found on the equator, where each cell has an angular area of $1.9\cdot 10^{-5}\mathrm{rad}^{2}$. We target a HEALPix grid with $n_{\mathrm{side}}=64$, where all of the $12\cdot 64^{2}$ pixels cover an angular area of $2.6\cdot 10^{-4}$, so that the ERA5 data has a higher spatial resolution everywhere on the globe and saturates the HEALPix pixel density.
@@ 4 PEAR: Pangu Equal Area
The architecture is constructed using a combination of 5 main layer types: patch embedding, windowed attention with alternating shifting, downsampling, upsampling and patch recovery. See Figure 3 for a schematic overview of the architecture, and Appendix A Table A1 for details on the layer parameters.
@@ 4.1 Patch embedding
The initial patch embedding uses a 1d convolution with kernel size 16 and stride 16 for the surface variables, and a 2d convolution with kernel size and stride $(16,2)$ for the upper variables. This corresponds to a patch size of $4\times 4$ in an equiangular grid. Both convolutions output 48 channels. The patch embedded tensors have shape $(\nicefrac{{3}}{{4}}n_{\mathrm{nside}}^{2},48)$ and $(\nicefrac{{3}}{{4}}n_{\mathrm{nside}}^{2},7,48)$. At this point the patch embedded surface variables are concatenated to the patch embedded upper variables, resulting in a single tensor of shape $(\nicefrac{{3}}{{4}}n_{\mathrm{nside}}^{2},8,48)$.
@@ 4.2 Windowed attention
We partition the tensors into windows using the nested structure of the HEALPix grid. The vertical direction is also partitioned into neighbouring levels. In terms of the latent tensor shape $(\nicefrac{{3}}{{4}}n_{\mathrm{nside}}^{2},8,48)$, partitioning using a window size of $(W_{\mathrm{hp}},W_{d})$ gives a windowed tensor of shape $(\frac{\nicefrac{{3}}{{4}}n_{\mathrm{nside}}^{2}}{W_{\mathrm{hp}}}\frac{8}{W_{\mathrm{d}}},W_{\mathrm{hp}}W_{d},48)$. Attention is now performed over the embeddings of the $W_{\mathrm{hp}}W_{d}$ voxels in each window. Note that the nested index structure of the HEALPix grid naturally supports the window partitioning in terms of contiguous memory. See Figure A1 for an illustration of the patches and first level windows corresponding to the tensor structure at the first attention layer.
In contrast to Pangu [bi2023], we use a simplified learned relative positional embedding. Because of the equal area grid cells in the HEALPix grid we can share relative positional embeddings between windows, and thus simply learn a relative positional embedding tensor $B$ of shape $(1,N_{\mathrm{heads}},\left(W_{d}W_{\mathrm{hp}}\right)^{2})$. The final attention computation is thus $\mathrm{Att}(Q,K,V)=\mathrm{SoftMax}\left(\frac{QK^{\top}}{\sqrt{d}}+B\right)V$, where $d$ is the embedding dimension. The simplified positional embedding accounts for most of the parameter savings compared to Pangu in Table 1.
@@ 5 Experiments
We evaluate PEAR against Pangu [zhuoqun2025] using the ERA5-lite subset described in Section 3.3. All experiments are carried out using $n_{\mathrm{side}}=64$ to saturate the angular resolution of the HEALPix grid using the available ERA5 data on the equiangular grid. We train all models using L1 loss, with the loss contribution from the surface variables weighted by $\frac{1}{4}$. We use the AdamW optimizer with weight decay $3\times 10^{-6}$ and learning rate $5\times 10^{-4}$. On a single A100, PEAR converges in 20 hours and Pangu in 40 hours.
To evaluate the medium-term forecasting ability of PEAR, we perform iterated model inference up to 10 times, resulting in forward time predictions of up to 10 days. At each lead time we calculate the average RMSE and anomaly correlation coefficient (ACC) [van2017probability] of all variables over the globe. See Appendix A for details.
To evaluate baseline-predictions on an equiangular grid, we apply the latitude weighting used in prior work [bi2023]. The equal area grid cells of HEALPix make this reweighting redundant for PEAR. The ACC measures the correlation between deviations from the climatology mean of predicted and ground truth forecasts, with a value of 1 indicating perfect agreement [hersbach2020era5]. The climatology average is subtracted to factor out seasonal variations that would otherwise improve the raw correlation between predictions and ground truth data. To calculate the climatology mean, we average each day of the year over the 11 years in the ERA5-lite dataset for every variable.
Pangu and Pangu-Large provide baselines against the same type of architecture where an equiangular grid is used for input, latent and output tensors. Since the overall architectural blocks are the same, we can match the depth, number of attention heads and embedding dimensions for each block. Pangu and PEAR share the same architecture hyperparameters, whereas Pangu-Large has 4 times larger embedding dimension. For Pangu and Pangu-Large we use a equiangular grid with $n_{\mathrm{lon}}=314$ and $n_{\mathrm{lat}}=157$, resulting in a relative difference of the total number of pixels of $0.3\%$, or 146 pixels.
@@
Table 2 lists the average ACC and RMSE for three different prediction lead times. PEAR consistently outperforms Pangu at longer lead times, and is often better than PanguLarge with 4 times larger embedding dimension. Additional evaluations are provided in the appendix.